How many grams of salt must be added to 90 grams of a 20% salt solution to make a 25% salt solution?
Answer by stanbon(75887)

How many grams of salt must be added to 90 grams of a 20% salt solution to make a 25% salt solution?
------
Equation:
salt + salt = salt
1*x + 0.20*90 = 0.25(x+90)
------
100x + 20*90 = 25x + 25*90
-----
75x = 5*90
-----
x = (1/15)90
x = 6 grams (amt of pure salt to add)
---------------------------------------
Cheers,
Stan H.